# Functional Distribution Networks (FDN)

### v2

@@ 6 Conclusion
We introduced Functional Distribution Networks (FDN), which amortize input-conditioned distributions over weights to produce predictive densities that remain sharp in-distribution yet expand appropriately under shift. Trained with a Monte Carlo objective and a $\beta$-weighted KL to a simple prior, FDN delivers strong rank calibration across tasks and near-ideal scale calibration on smooth and piecewise-smooth shifts (step, quadratic), as evidenced by Spearman correlation, MSE–Var slope/intercept, $\Delta\mathrm{Var}$, $\Delta\mathrm{CRPS}$, and $\mathrm{AURC}$. Under a fair protocol that matches parameters, updates, and predictive-sample budgets, FDN remains competitive with standard Bayesian, ensemble, dropout, and hypernetwork baselines on both 1D toy functions and real UCI-style regression benchmarks, and provides uncertainty that is practically useful for abstention and risk-aware inference.
Looking forward, we see several promising directions. On the modeling side, stronger variance scaling mechanisms (temperature/floors, layer-wise $\beta$ scheduling, structured priors) and frequency-aware conditioning may close the remaining gap on highly oscillatory OOD tasks. On the systems side, adapter-style deployments that apply FDN only to a small subset of middle or head layers could allow one to inject uncertainty awareness in new or existing large-scale architectures with modest overhead. Finally, extending FDN to classification, sequence models, and structured prediction—as well as integrating it with other uncertainty-aware modules (e.g., Neural Processes or diffusion-style generative priors over weights)—could yield a general toolkit for calibrated, shift-aware deep learning in practical applications.
@@ Appendix A Unified view via $q_{\phi}(\theta\mid x)$
$\beta\sum_{\ell=1}^{L}\displaystyle D_{\mathrm{KL}}(q_{\phi}(\theta_{\ell}\mid x)\|p_{0}(\theta_{\ell})).$
@@ A Stochastic Hypernetwork outputs a global posterior (or context-only):
and inference averages predictions over sampled masks (gal2016dropout).
@@ Appendix B Training objective
$\log p(y\mid x)\ \geq\ \underbrace{\mathbb{E}_{q_{\phi}(\theta\mid x)}\big[\log p(y\mid x,\theta)\big]}_{\text{data term}}\ -\ \underbrace{D_{\mathrm{KL}}\!\big(q_{\phi}(\theta\mid x)\,\|\,p_{0}(\theta)\big)}_{\text{regularizer}},$ (1)
@@
$\mathcal{L}_{\beta\text{-ELBO}}=-\frac{1}{K}\sum_{k=1}^{K}\log p\!\big(y\mid x,\theta^{(k)}\big)\;+\;\beta\,D_{\mathrm{KL}}\!\big(q_{\phi}(\theta\mid x)\,\|\,p_{0}(\theta)\big),\qquad\theta^{(k)}\!\sim q_{\phi}(\theta\mid x).$ (2)
Here $\beta{=}1$ recovers standard VI; $\beta\!\neq\!1$ implements capacity control / tempered VI (higgins2017beta; alemi2017deep; dziugaite2017computing). We use simple warm-ups for $\beta$ early in training.
@@
which implicitly accounts for the KL via the weights and typically needs no extra $\beta$ (burda2016importance). We report main results with (A) for simplicity and stability.
@@
$D_{\mathrm{KL}}\big(q_{\phi,\ell}(\theta_{\ell}\!\mid a_{\ell-1}^{(k)}(x))\ \|\ p_{0}\big)\quad\text{with}\quad\mathbb{E}_{k}[\cdot]\ \text{across samples }k.$
@@
$D_{\mathrm{KL}}\!\left(\mathcal{N}(\mu,\mathrm{diag}\,\sigma^{2})\,\|\,\mathcal{N}(0,\sigma_{0}^{2}I)\right)=\tfrac{1}{2}\sum_{j}\!\left(\frac{\sigma_{j}^{2}+\mu_{j}^{2}}{\sigma_{0}^{2}}-1-\log\frac{\sigma_{j}^{2}}{\sigma_{0}^{2}}\right),$
@@ Appendix D Heteroscedastic likelihood and variance decomposition
$\mathcal{L}_{\text{Gauss}}^{(i)}$ $=\frac{1}{2K}\sum_{k=1}^{K}\Big[\big\|y_{i}-\mu_{\theta_{k}}(x_{i})\big\|^{2}_{\left(\Sigma_{\theta_{k}}(x_{i})\right)^{-1}}\;+\;\log\det\!\big(2\pi\,\Sigma_{\theta_{k}}(x_{i})\big)\Big]\;+\;\beta\,D_{\mathrm{KL}}\!\big(q_{\phi}(\theta\mid x_{i})\,\|\,p_{0}(\theta)\big),$ (4)
@@
$\log p(y_{i}\mid x_{i})\ \geq\ \underbrace{\mathbb{E}_{q_{\phi}}\!\big[\log p(y_{i}\mid x_{i},\theta)\big]}_{\text{data term}}\;-\;\underbrace{D_{\mathrm{KL}}\!\big(q_{\phi}(\theta\mid x_{i})\,\|\,p_{0}(\theta)\big)}_{\text{regularizer}}.$
@@
$\mathcal{L}_{\text{ELBO}}^{(i)}=\frac{1}{2\sigma^{2}}\ \mathbb{E}_{q_{\phi}(\theta\mid x_{i})}\!\big[(y_{i}-f_{\theta}(x_{i}))^{2}\big]\;+\;D_{\mathrm{KL}}\!\big(q_{\phi}(\theta\mid x_{i})\,\|\,p_{0}(\theta)\big)\;+\;\tfrac{1}{2}\log(2\pi\sigma^{2}).$
@@
$\boxed{\;\mathcal{L}_{\text{ELBO}}^{(i)}\approx\frac{1}{2K\sigma^{2}}\sum_{k=1}^{K}\big(y_{i}-f_{\theta^{(k)}}(x_{i})\big)^{2}\;+\;D_{\mathrm{KL}}\!\big(q_{\phi}(\theta\mid x_{i})\,\|\,p_{0}(\theta)\big)\;+\;\tfrac{1}{2}\log(2\pi\sigma^{2})\;}.$
@@
$(2\sigma^{2})\,\mathcal{L}_{\text{ELBO}}^{(i)}\doteq\frac{1}{K}\sum_{k=1}^{K}\big(y_{i}-f_{\theta^{(k)}}(x_{i})\big)^{2}\;+\;\underbrace{(2\sigma^{2})}_{\beta}\,D_{\mathrm{KL}}\!\big(q_{\phi}(\theta\mid x_{i})\,\|\,p_{0}(\theta)\big),$
@@
$\mathcal{L}_{\text{het}}^{(i)}=\frac{1}{2K}\sum_{k=1}^{K}\Bigg[\frac{\big(y_{i}-f_{\theta^{(k)}}(x_{i})\big)^{2}}{\sigma_{\theta^{(k)}}^{2}(x_{i})}\;+\;\log\!\big(2\pi\,\sigma_{\theta^{(k)}}^{2}(x_{i})\big)\Bigg]\;+\;\beta\,D_{\mathrm{KL}}\!\big(q_{\phi}(\theta\mid x_{i})\,\|\,p_{0}(\theta)\big),$
@@
and obtain the total predictive variance via the decomposition equation 6,
@@
For $d_{y}>1$, replace squared deviations by outer products to estimate covariances, in accordance with equation 7.
@@ Appendix E FDN Training and Prediction Algorithm
11:     $\Sigma_{\mathrm{KL}}\mathrel{+}=\displaystyle D_{\mathrm{KL}}\big(q_{\phi}^{l}(\theta_{l}\!\mid\!c_{l})\,\|\,p_{0}^{l}(\theta_{l})\big)$
@@ Appendix G Computational complexity and scalability
In practice, Hypernetworks need not be applied to every layer of a deep backbone. FDN is layer-local, so one can restrict stochastic, input-conditioned weights to a small subset of higher layers (or even just the predictive head), keeping earlier blocks deterministic and frozen. This is analogous in spirit to LoRA and adapter modules (hu2022lora; houlsby2019parameter): a narrow, trainable “uncertainty adapter” is inserted on top of a largely fixed backbone, so the additional parameters and compute scale with the adapter width rather than with the full network depth or width. In such configurations the overall parameter and FLOP overhead of FDN remains a small fraction of the backbone, even for large architectures, while still enabling input-dependent uncertainty where it is most needed.
@@ Appendix H Evaluation metrics and calibration diagnostics
LLM disclosure
ChatGPT assisted with minor copy-editing, LaTeX phrasing, and bibliography chores (suggesting candidate references and drafting BibTeX). The authors independently reviewed the literature and verified all citation metadata (titles, authors, venues, DOIs/arXiv).

### v3

@@ 6 Conclusion
We introduced Functional Distribution Networks (FDN), which amortize input-conditioned distributions over weights to produce predictive densities that remain sharp in-distribution and expand under shift. Trained with a Monte Carlo objective and a $\beta$-weighted KL prior, FDN achieves strong rank calibration across tasks and near-ideal scale calibration on smooth and piecewise-smooth shifts, as measured by Spearman correlation, MSE–Var fit, $\Delta\mathrm{Var}$, $\Delta\mathrm{CRPS}$, and AURC. Under matched parameter, update, and sampling budgets, FDN is competitive with Bayesian, ensemble, dropout, and hypernetwork baselines on both 1D toy functions and UCI-style regression benchmarks, while providing uncertainty suitable for abstention and risk-aware inference.
Looking ahead, several directions are promising. On the modeling side, stronger variance-scaling mechanisms (e.g., temperature or layer-wise $\beta$ scheduling, structured priors) and frequency-aware conditioning may reduce under-scaling on highly oscillatory OOD tasks. On the systems side, adapter-style deployments that apply FDN to a small subset of layers could inject uncertainty awareness into large-scale models with modest overhead. Finally, extending FDN to classification, sequence models, and structured prediction, or integrating it with other uncertainty-aware modules (e.g., Neural Processes or diffusion-style priors over weights), may enable a general toolkit for calibrated, shift-aware deep learning.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Unified view via $q_{\phi}(\theta\mid x)$
$\beta\sum_{\ell=1}^{L}\displaystyle KL(q_{\phi}(\theta_{\ell}\mid x)\|p_{0}(\theta_{\ell})).$
@@ A Stochastic Hypernetwork outputs a global posterior (or context-only):
and inference averages predictions over sampled masks (Gal and Ghahramani, 2016).
@@ Appendix B Training objective
$\log p(y\mid x)\ \geq\ \underbrace{\mathbb{E}_{q_{\phi}(\theta\mid x)}\big[\log p(y\mid x,\theta)\big]}_{\text{data term}}\ -\ \underbrace{KL\!\big(q_{\phi}(\theta\mid x)\,\|\,p_{0}(\theta)\big)}_{\text{regularizer}},$ (1)
@@
$\mathcal{L}_{\beta\text{-ELBO}}=-\frac{1}{K}\sum_{k=1}^{K}\log p\!\big(y\mid x,\theta^{(k)}\big)\;+\;\beta\,KL\!\big(q_{\phi}(\theta\mid x)\,\|\,p_{0}(\theta)\big),\qquad\theta^{(k)}\!\sim q_{\phi}(\theta\mid x).$ (2)
Here $\beta{=}1$ recovers standard VI; $\beta\!\neq\!1$ implements capacity control / tempered VI (Higgins and others, 2017; Alemi et al., 2017; Dziugaite and Roy, 2017). We use simple warm-ups for $\beta$ early in training.
@@
which implicitly accounts for the KL via the weights and typically needs no extra $\beta$ (Burda et al., 2016). We report main results with (A) for simplicity and stability.
@@
$KL\big(q_{\phi,\ell}(\theta_{\ell}\!\mid a_{\ell-1}^{(k)}(x))\ \|\ p_{0}\big)\quad\text{with}\quad\mathbb{E}_{k}[\cdot]\ \text{across samples }k.$
@@
$KL\!\left(\mathcal{N}(\mu,\mathrm{diag}\,\sigma^{2})\,\|\,\mathcal{N}(0,\sigma_{0}^{2}I)\right)=\tfrac{1}{2}\sum_{j}\!\left(\frac{\sigma_{j}^{2}+\mu_{j}^{2}}{\sigma_{0}^{2}}-1-\log\frac{\sigma_{j}^{2}}{\sigma_{0}^{2}}\right),$
@@ Appendix D Heteroscedastic likelihood and variance decomposition
$\mathcal{L}_{\text{Gauss}}^{(i)}$ $=\frac{1}{2K}\sum_{k=1}^{K}\Big[\big\|y_{i}-\mu_{\theta_{k}}(x_{i})\big\|^{2}_{\left(\Sigma_{\theta_{k}}(x_{i})\right)^{-1}}\;+\;\log\det\!\big(2\pi\,\Sigma_{\theta_{k}}(x_{i})\big)\Big]\;+\;\beta\,KL\!\big(q_{\phi}(\theta\mid x_{i})\,\|\,p_{0}(\theta)\big),$ (4)
@@
$\log p(y_{i}\mid x_{i})\ \geq\ \underbrace{\mathbb{E}_{q_{\phi}}\!\big[\log p(y_{i}\mid x_{i},\theta)\big]}_{\text{data term}}\;-\;\underbrace{KL\!\big(q_{\phi}(\theta\mid x_{i})\,\|\,p_{0}(\theta)\big)}_{\text{regularizer}}.$
@@
$\mathcal{L}_{\text{ELBO}}^{(i)}=\frac{1}{2\sigma^{2}}\ \mathbb{E}_{q_{\phi}(\theta\mid x_{i})}\!\big[(y_{i}-f_{\theta}(x_{i}))^{2}\big]\;+\;KL\!\big(q_{\phi}(\theta\mid x_{i})\,\|\,p_{0}(\theta)\big)\;+\;\tfrac{1}{2}\log(2\pi\sigma^{2}).$
@@
$\boxed{\;\mathcal{L}_{\text{ELBO}}^{(i)}\approx\frac{1}{2K\sigma^{2}}\sum_{k=1}^{K}\big(y_{i}-f_{\theta^{(k)}}(x_{i})\big)^{2}\;+\;KL\!\big(q_{\phi}(\theta\mid x_{i})\,\|\,p_{0}(\theta)\big)\;+\;\tfrac{1}{2}\log(2\pi\sigma^{2})\;}.$
@@
$(2\sigma^{2})\,\mathcal{L}_{\text{ELBO}}^{(i)}\doteq\frac{1}{K}\sum_{k=1}^{K}\big(y_{i}-f_{\theta^{(k)}}(x_{i})\big)^{2}\;+\;\underbrace{(2\sigma^{2})}_{\beta}\,KL\!\big(q_{\phi}(\theta\mid x_{i})\,\|\,p_{0}(\theta)\big),$
@@
$\mathcal{L}_{\text{het}}^{(i)}=\frac{1}{2K}\sum_{k=1}^{K}\Bigg[\frac{\big(y_{i}-f_{\theta^{(k)}}(x_{i})\big)^{2}}{\sigma_{\theta^{(k)}}^{2}(x_{i})}\;+\;\log\!\big(2\pi\,\sigma_{\theta^{(k)}}^{2}(x_{i})\big)\Bigg]\;+\;\beta\,KL\!\big(q_{\phi}(\theta\mid x_{i})\,\|\,p_{0}(\theta)\big),$
@@
and obtain the total predictive variance via the decomposition (6),
@@
For $d_{y}>1$, replace squared deviations by outer products to estimate covariances, in accordance with (7).
@@ Appendix E FDN Training and Prediction Algorithm
11:     $\Sigma_{\mathrm{KL}}\mathrel{+}=\displaystyle KL\big(q_{\phi}^{l}(\theta_{l}\!\mid\!c_{l})\,\|\,p_{0}^{l}(\theta_{l})\big)$
@@ Appendix G Computational complexity and scalability
In practice, Hypernetworks need not be applied to every layer of a deep backbone. FDN is layer-local, so one can restrict stochastic, input-conditioned weights to a small subset of higher layers (or even just the predictive head), keeping earlier blocks deterministic and frozen. This is analogous in spirit to LoRA and adapter modules (Hu et al., 2022; Houlsby et al., 2019): a narrow, trainable “uncertainty adapter” is inserted on top of a largely fixed backbone, so the additional parameters and compute scale with the adapter width rather than with the full network depth or width. In such configurations the overall parameter and FLOP overhead of FDN remains a small fraction of the backbone, even for large architectures, while still enabling input-dependent uncertainty where it is most needed.